OpenStudy (anonymous):

find the odd one out: 1 4 16 27 256 3125

OpenStudy (anonymous):

Seems like 16. 1=1^1 4=2^2 27=3^3 256=4^4 3125=5^5

OpenStudy (anonymous):

wow thats gr8....i toltally agree wid u..